# On the principle of linearized stability for quasilinear evolution equations in time-weighted spaces

### v1

@@ 1. Introduction
$\|u\|_{C_{\mu}((0,t],E_{\xi})}\leq\frac{L}{2}\qquad\text{and}\qquad\|u\|_{C([0,t],E_{\alpha})}\leq\frac{R}{2}$
for each $t\in(0,t_{1})$, hence $t_{1}=t^{+}(v^{0})$ and, together with (1.11e) and (1.14), we conclude, similarly as in the proof of Theorem 1.3, that $t^{+}(v^{0})=\infty$.
Define now
@@
and together with (1.15d), (3.1), and the latter estimate we deduce, analogously to (3.5)-(3.6), that
@@ 4. Applications
In this section we apply our theory to various semilinear and quasilinear evolution problems of parabolic type, examining both critical and non-critical regimes, to explore the stability properties of their equilibrium solutions. The applications include quasilinear problems with quadratic semilinearities (see Example 4.1), a parabolic-parabolic chemotaxis system (Example 4.2), and a quasilinear evolution equation in critical spaces exhibiting scaling invariance (Example 4.3).
@@ 4.1. Quaslilinear Problems with Quadratic Semilinearity
$0<\gamma=\gamma_{*}<\beta<\xi<1\qquad\text{and}\qquad 2\xi-1-\gamma\leq\alpha\in(\beta,\xi)\,,$ (4.2c)
noticing that $\alpha_{\rm crit}=\alpha_{\rm crit}^{*}=2\xi-1-\gamma$. Then, $f:E_{\xi}\to E_{\gamma}$ with $f(u):=Q(u,u)$ satisfies (1.4) and (1.8d), both with $q=q_{*}=2$.
@@
Of course, as pointed out in Remark 1.1, also other interpolation functors than the complex functor $[\cdot,\cdot]_{\theta}$ (such as real or continuous interpolation functors) can be considered.
In the critical case $2\xi-1-\gamma=\alpha\in(\gamma,\xi)$, Theorem 4.1 has been established previously in the particular context of a semilinear asymptotic model for atmospheric flows describing morning glory clouds, see [12, Theorem 4.1]. For similar results for semilinear parabolic problems with general superlinear nonlinearites we refer to [12, Corollary 1.4] and [12, Example 4.2]. We also refer to [17, Corollary 2.2] for a related exponential stability result in the semilinear case within the framework of maximal $L_{p}$-regularity.
@@ 4.2. A Parabolic-Parabolic Chemotaxis System
It is well-known that for sufficiently smooth non-negative initial data (4.3) possesses a unique bounded global classical solution provided that $\Omega$ is convex and $\kappa$ is sufficiently large, see e.g. [22]. We prove herein that (4.3) is locally well-posed for more general initial data and that the equilibrium solution ${(u_{1},v_{1})=(0,0)}$ is always unstable. Moreover, under certain restrictions of the coefficients $\chi$ and $\kappa$, we establish the exponential stability of the equilibrium $(u_{2},v_{2})=(1,1)$.
Before presenting our result in Theorem 4.2, we define for $p\in(1,\infty)$ the Banach spaces
@@
$\big{\{}(F_{\theta},B_{\theta})\,:\,-1\leq\theta<\infty\big{\}}$
be the interpolation-extrapolation scale generated by $(F_{0},B_{0})$ and the complex interpolation functor $[\cdot,\cdot]_{\theta}$ (see [3, §6] and [4, §V.1]). Then,
@@
$F_{\theta}\doteq H^{2\theta}_{p,N}(\Omega):=\left\{\begin{array}[]{ll}\{v\in H_{p}^{2\theta}(\Omega)\,:\,\partial_{\nu}v=0\text{ on }\partial\Omega\}\,,&1+\frac{1}{p}<2\theta<3+\frac{1}{p}\,,\\[3.0pt] H_{p}^{2\theta}(\Omega)\,,&-1+\frac{1}{p}<2\theta<1+\frac{1}{p}\,,\end{array}\right.$ (4.5)
see [3, Theorem 7.1; Equation (7.5)] [^3] This property is stated in [3] for $-1+\frac{1}{p}<2\theta\leq 2$. However, since $(1-\Delta_{N})^{-1}\in\mathcal{L}(H_{p}^{2\theta-2}(\Omega),H_{p}^{2\theta}(\Omega))$ for $2<2\theta<3+1/p$, see [21, Theorem 5.5.1], we obtain the full range in (4.5). Moreover, since $\Delta_{N}-1$ has bounded imaginary powers, see e.g. from [4, III. Examples 4.7.3 (d)], we infer from [3, Remarks 6.1 (d)] that
$[F_{\beta},F_{\alpha}]_{\theta}\doteq F_{(1-\theta)\beta+\theta\alpha}\,,\qquad-1\leq\beta<\alpha\,,\quad\theta\in(0,1)\,.$ (4.6)
@@
[…]
@@
$q=q_{*}=2\qquad\text{and}\qquad 0<\gamma=\gamma_{*}:=\frac{\varepsilon}{3}<\alpha:=\varepsilon<\xi=\frac{1+\varepsilon}{2}<1\,,$
@@
$\alpha_{\rm crit}=\alpha_{\rm crit}^{*}=2\xi-1-\gamma=\frac{2\varepsilon}{3}<\alpha\in(\gamma,\xi)\,,$
@@
[…]
@@
[…]
hence (4.8) (with $\bar{w}=0$) implies that
$\|f(w+w_{2})-f(w_{2})-\partial f(w_{2})w\|_{\gamma}=\|f(w)\|_{\gamma}\leq C\|w\|_{\xi}^{2},\qquad w\in E_{\xi}\,.$
Thus, the assumptions (1.15a)-(1.15d) are all satisfied within this framework.
It remains to verify that (1.15e) is also satisfied in the context of (4.3). Let thus $\lambda={\rm Re\,}\lambda+i\,{\rm Im\,}\lambda\in\mathbb{C}$ be an eigenvalue of ${A+\partial f(w_{2})}$ with eigenvector $0\neq w=(u,v)\in H^{2}_{N}(\overline{\Omega})^{2}$ by elliptic regularity, see e.g. [1, Theorem 15.2]. It then holds
@@
hence ${\rm Re\,}\lambda<0$, which proves (1.15e). Assertion (iii) is now a direct consequence of Theorem 1.6. ∎
@@ 4.3. A Quasilinear Problem with Scaling Invariance
$\partial_{t}u={\rm div}(a(u)\nabla u)+|\nabla u|^{\kappa}\qquad\text{in $\Omega$\,,\, $t>0$\,,}$ (4.11a)
@@
can be identified (via a scaling invariance argument) as a critical space for (4.11). The next results establishes the exponential stability of the zero solution to (4.11) in this critical space.
@@
$\big{\{}(F_{\theta},B_{\theta})\,:\,-1\leq\theta<\infty\big{\}}$
be the interpolation-extrapolation scale generated by $(F_{0},B_{0})$ and the complex interpolation functor $[\cdot,\cdot]_{\theta}$ (see [3, §6] and [4, §V.1]), that is,
@@
$[F_{\beta},F_{\alpha}]_{\theta}\doteq F_{(1-\theta)\beta+\theta\alpha}\,,\qquad-1\leq\beta<\alpha\,,\quad\theta\in(0,1)\,.$ (4.15)
@@
$0<\gamma=\gamma_{*}:=\tau<\beta:=\tau+\frac{\bar{s}}{2}<\alpha:=\alpha_{\rm crit}<\xi:=\tau+\frac{s}{2}<1\,,$ (4.16)
@@
$\alpha_{\rm crit}=\alpha_{\rm crit}^{*}:=\frac{\kappa\xi-1-\gamma}{\kappa-1}=\tau+\frac{s_{c}}{2}\in(\beta,\xi)\,.$
@@
We next verify (1.8) for the stationary solution $v_{*}:=0\in E_{1}$. To this end we first infer from (4.18) that the map $f:E_{\xi}\to E_{\gamma}$ is Fréchet differentiable in 00 with $\partial f(0)=0$ and
$\|f(w)\|_{\gamma}\leq\kappa\|w\|_{\xi}^{\kappa}\,,\qquad w\in E_{\xi}\,.$
Since the spectral bound $s(\Delta_{D})=\sup\{\mathrm{Re}\,\lambda:\lambda\in\sigma(\Delta_{D})\}$ is negative and $v_{*}=0$ (in which case the left hand side of (1.8e) vanishes), assumptions (1.8) are satisfied. We are thus in a position to apply Theorem 1.3 and establish in this way the exponential stability of the zero solution to (4.17) in $H^{s_{c}}_{p,D}(\Omega)$, see (4.12). ∎

### v2

@@ 1. Introduction
$\|u\|_{C_{\mu}((0,t],E_{\xi})}\leq\frac{L}{2}\qquad\text{and}\qquad\|u\|_{C([0,t],E_{\alpha})}\leq\frac{R}{2}$ (3.12)
for each $t\in(0,t_{1})$, hence $t_{1}=t^{+}(v^{0})$. In particular, the solution $v(\cdot;v^{0}):[0,t^{+}(v^{0}))\to E_{\alpha}$} to (1.10) satisfies, in view of (3.12),
$\liminf_{t\nearrow t^{+}(v^{0})}\mathrm{dist}_{E_{\alpha}}\big(v(t;v^{0}),\partial O_{\alpha}\big)>0$
and
$\|v(t;v^{0})\|_{\xi}\leq\|v_{*}\|_{\xi}+Lt^{-\mu}\,,\qquad t\in(0,t^{+}(v^{0}))\,.$
This property together with (1.11e) ensures that
$\underset{t\nearrow t^{+}(v^{0})}{\limsup}\|f(v(t;v^{0}))\|_{0}<\infty\,$,
and Theorem 1.5 implies now that $t^{+}(v^{0})=\infty$.
Define
@@
and together with (1.16b), (3.5), and the latter estimate we deduce, analogously to (3.10)-(3.11), that
@@ 4. Applications
In this section we apply our theory to various semilinear and quasilinear evolution problems of parabolic type, examining both critical and noncritical regimes, to explore the stability properties of their equilibrium solutions. The applications include quasilinear problems with quadratic semilinearities (see Example 4.1), a parabolic-parabolic chemotaxis system (Example 4.2), and a quasilinear evolution equation in critical spaces exhibiting scaling invariance (Example 4.3).
@@ 4.1. Quaslilinear Problems with Quadratic Semilinearity
$0<\gamma<\beta<\xi<1\,,\qquad q:=2\,,\qquad 2\xi-1-\gamma\leq\alpha\in(\beta,\xi)\,.$ (4.2c)
In the semilinear case when $A(u)=A\in\mathcal{H}(E_{1},E_{0})$, we put $\gamma_{*}:=\gamma$ and $q_{*}:=q=2$, noticing that thus $\alpha_{\rm crit}=\alpha_{\rm crit}^{*}=2\xi-1-\gamma$. Then, $f:E_{\xi}\to E_{\gamma}$ with $f(u):=Q(u,u)$ satisfies (1.4) and (1.16b).
@@
Of course, as pointed out in Remark 1.1, also other interpolation functors than the complex functor $\cdot,\cdot]_{\theta}$ (such as real or continuous interpolation functors) can be considered.
In the critical case $2\xi-1-\gamma=\alpha\in(\gamma,\xi)$, Theorem 4.1 has been established previously in the particular context of a semilinear asymptotic model for atmospheric flows describing morning glory clouds, see [12, Theorem 4.1]. It is worth noting that Theorem 4.1 yields an exponential stability result which, to the best of our knowledge, cannot be obtained via the classical approach [13, 3, 7]. For similar results for semilinear parabolic problems with general superlinear nonlinearities we refer to [12, Corollary 1.4] and [12, Example 4.2]. We also refer to [17, Corollary 2.2] for a related exponential stability result in the semilinear case within the framework of maximal $L_{p}$-regularity.
@@ 4.2. A Parabolic-Parabolic Chemotaxis System
It is well-known that for sufficiently smooth non-negative initial data, (4.3) possesses a unique bounded global classical solution provided that $\Omega$ is convex and $\kappa$ is sufficiently large, see e.g. [22]. We prove herein that (4.3) is locally well-posed for more general initial data and that the equilibrium solution $(u_{1},v_{1})=(0,0)$} is always unstable. Moreover, under certain restrictions on the coefficients $\chi$ and $\kappa$, we establish the exponential stability of the equilibrium $(u_{2},v_{2})=(1,1)$. The advantages of the parabolic theory based on time-weighted function spaces in the context of (4.3) are highlighted in [14], where local well-posedness is established for a more general class of models, assuming initial data $u_{0}\in L_{2}(\Omega)$ in the physically relevant dimensions $n\leq 3$}. Moreover, it is shown in [14] that global $L_{2}$-bounds on $u$ suffice to guarantee the global existence of solutions to (4.3). In this context, our result in Theorem 4.2 complements the analysis in [14] by providing a corresponding stability result within the same low-regularity functional framework.
Before presenting Theorem 4.2, we define for $p\in(1,\infty)$ the Banach spaces
@@
$\big\{(F_{\theta},B_{\theta})\,:\,-1\leq\theta<\infty\big\}\$
be the interpolation-extrapolation scale generated by $(F_{0},B_{0})$ and the complex interpolation functor $\cdot,\cdot]_{\theta}$ (see [3, §6] and [4, §V.1]). Then,
@@
[…]
see [3, Theorem 7.1; Equation (7.5)] [^2] This property is stated in [3] for $1+\frac{1}{p}<2\theta\leq 2$. However, since $(1-\Delta_{N})^{-1}\in\mathcal{L}(H_{p}^{2\theta-2}(\Omega),H_{p}^{2\theta}(\Omega))$ for $2<2\theta<3+1/p$, see [21, Theorem 5.5.1], we obtain the full range in (4.5).. Moreover, since $\Delta_{N}-1$ has bounded imaginary powers, see e.g. from [4, III. Examples 4.7.3 (d)], we infer from [3, Remarks 6.1 (d)] that
$F_{\beta},F_{\alpha}]_{\theta}\doteq F_{(1-\theta)\beta+\theta\alpha}\,,\qquad-1\leq\beta<\alpha\,,\quad\theta\in(0,1)\,.$ (4.6)
@@
$E_{\theta}=H_{p,N}^{2\theta-2\varepsilon}(\Omega)\times H_{q,N}^{1+2\theta-2\varepsilon}(\Omega)\,,\qquad 2\theta\in[0,2]\setminus\{2\varepsilon+1+1/p\,,\,2\varepsilon+1/2p\}\,.$
@@
$q:=2\qquad\text{and}\qquad 0<\gamma:=\frac{\varepsilon}{3}<\alpha:=\varepsilon<\xi:=\frac{1+\varepsilon}{2}<1\,$,
@@
$\alpha_{\rm crit}=2\xi-1-\gamma=\frac{2\varepsilon}{3}<\alpha\in(\gamma,\xi)\,$,
@@
$\partial f(\bar{w})w=-\big(\chi\mathrm{div}\big(u\nabla\bar{v}+\bar{u}\nabla v)+2\kappa u\bar{u}\,,\,0\big)\,,\qquad w=(u,v)\,,\,\bar{w}=(\bar{u},\bar{v})\in E_{\xi}\,.$
@@
$\partial f(w_{2})w=\big(\!-\chi\Delta v-2\kappa u\,,\,0\big)\,,\qquad w=(u,v)\in E_{\xi}\,$,
and, in order to apply Theorem 1.6, it remains to verify that (1.15c) is satisfied in the context of (4.3). Let thus $\lambda={\rm Re\,}\lambda+i\,{\rm Im\,}\lambda\in\mathbb{C}$ be an eigenvalue of $A+\partial f(w_{2})$} with eigenvector $0\neq w=(u,v)\in H^{2}_{N}(\Omega)^{2}$ by elliptic regularity, see e.g. [1, Theorem 15.2]. It then holds
@@
hence ${\rm Re\,}\lambda<0$, which proves (1.15c). Assertion (iii) is now a direct consequence of Theorem 1.6. ∎
@@ 4.3. A Quasilinear Problem with Scaling Invariance
[…]
@@
can be identified (via a scaling invariance argument) as a critical space for (4.11). The next result, which is new to the best of our knowledge, establishes the exponential stability of the zero solution to (4.11) in this critical space. For a related exponential stability result in a critical Besov space we refer to [17, Example 2].
@@
$\big\{(F_{\theta},B_{\theta})\,:\,-1\leq\theta<\infty\big\}\$
be the interpolation-extrapolation scale generated by $(F_{0},B_{0})$ and the complex interpolation functor $\cdot,\cdot]_{\theta}$ (see [3, §6] and [4, §V.1]), that is,
@@
$F_{\beta},F_{\alpha}]_{\theta}\doteq F_{(1-\theta)\beta+\theta\alpha}\,,\qquad-1\leq\beta<\alpha\,,\quad\theta\in(0,1)\,.$ (4.15)
@@
$0<\gamma:=\tau<\beta:=\tau+\frac{\bar{s}}{2}<\alpha:=\tau+\frac{s}{2}<1\,$, (4.16)
@@
$\alpha_{\rm crit}=\frac{\kappa\xi-1-\gamma}{\kappa-1}=\tau+\frac{s_{c}}{2}\in(\beta,\xi)\,.$
@@
We next verify (1.8) for the stationary solution $v_{*}:=0\in E_{1}$. To this end we first infer from (4.18) that the map $f:E_{\xi}\to E_{\gamma}$ is Fréchet differentiable in 0 with $\partial f(0)=0$. Since the spectral bound
$s(\Delta_{D})=\sup\{\mathrm{Re}\,\lambda:\lambda\in\sigma(\Delta_{D})\}$
is negative, assumptions (1.8) are satisfied. We are thus in a position to apply Theorem 1.3 and establish in this way the exponential stability of the zero solution to (4.17) in $H^{s_{c}}_{p,D}(\Omega)$, see (4.12). ∎
Acknowledgement
We express our gratitude to the referees for their careful reading of the manuscript and for their valuable suggestions, which have helped to improve the paper.